# Birational Geometry of Special Quotient Foliations and Chazy’s Equations

### v1

@@ 3.1. Brunella’s very special foliation, $\mathcal{F}_{3}$
It is tangent to the coordinate triangle $\Delta_{3}:XZY=0$, and has three reduced non-degenerate singular points at its vertices. It is preserved by the linear automorphism of order three
@@
  • first, a blow-up $\sigma$ of the node $p$ of the link $C$ transforms it into a curve $\overline{C}$ of self-intersection $-1$;
@@ 3.2. Santos’s foliation $\mathcal{F}_{4}$
for $\mathrm{i}=\sqrt{-1}$. It is tangent to the cycle $\Delta_{4}$ of four lines formed by $\mathbb{P}^{1}\times\{0\}$, $\mathbb{P}^{1}\times\{\infty\},$$\{0\}\times{\mathbb{P}}^{1}$, and {}×$\{\infty\}\times\mathbb{P}^{1}$, and has reduced non-degenerate singularities at its vertices. The order-four automorphism $T_{4}$ of $\mathbb{P}^{1}\times\mathbb{P}^{1}$
@@
permutes cyclically the four lines of $\Delta_{4}$, and preserves the foliation $\mathcal{E}_{4}$. It acts freely in a neighborhood of $\Delta_{4}$. The transformation $T_{4}$ has two fixed points, $(1:1,1:1)$ and $(-1:1,-1:1)$, at which the eigenvalues of the derivative of $T_{4}$ are $\mathrm{i}$ and $-\mathrm{i}$. It also has an orbit of length two, formed by $(-1:1,1:1)$ and $(1:1,-1:1)$, at which the derivative of $T^{2}$ has twice the eigenvalue $-1$. The variety $(\mathbb{P}^{1}\times\mathbb{P}^{1})/T_{4}$ has four singular points, two of type $A_{3}$ and one of type $A_{1}$. The minimal desingularization $M_{4}\to(\mathbb{P}^{1}\times\mathbb{P}^{1})/T_{4}$ is endowed with a foliation, Santos’s foliation $\mathcal{F}_{4}$, coming from $\mathcal{E}_{4}$, and a link $C$ (in the sense of Section 1), image of $\Delta_{4}$, with $C^{2}=2$ (see [San17]).
@@ 3.3. Santos’s foliation $\mathcal{F}_{6}$
Let us consider again the foliation $\mathcal{E}_{3}$ of $\mathbb{P}^{2}$ of Section 3.1 of Eq. (15). Let $\sigma:\mathrm{Bl}_{3}(\mathbb{P}^{2})\to\mathbb{P}^{2}$ be the composition of the blowing-ups at the vertices $(0:0:1)$, $(0:1:0)$, and $(1:0:0)$. Consider the cycle $\Delta_{6}$ of six $(-1)$-curves on $\mathrm{Bl}_{3}(\mathbb{P}^{2})$ composed by the transformed of the coordinate triangle of $\mathbb{P}^{2}$ by $\sigma$, and by the three exceptional lines of $Q$. Denote by $\mathcal{E}_{6}$ the transformed foliation of $\mathcal{E}_{3}$ by $Q$; it is tangent to $\Delta_{6}$, and has reduced non-degenerate singularities at its six vertices. The linear transformation $T_{3}$ of $\mathbb{P}^{2}$ of Eq. (16) commutes with the Cremona involution $Q$ in Eq. (17), and $T_{3}\circ Q$ is a birational automorphism of order six of $\mathbb{P}^{2}$ preserving $\mathcal{L}$. It induces a biholomorphism $T_{6}$ of $\mathrm{Bl}_{3}(\mathbb{P}^{2})$, which permutes cyclically the six rational curves of $\Delta_{6}$, and preserves the foliation $\mathcal{E}_{6}$. The automorphism $T_{6}$ has a fixed point coming from $(1:1:1)$, the common fixed point of $T_{3}$ and $Q$; an orbit of order two coming from the other fixed points of $T_{3}$, $(1:\omega:\omega^{2})$ and $(1:\omega^{2}:\omega)$; and an orbit of order three, formed by the other fixed points of $Q$, $(1:1:-1)$, $(-1:1:1)$ and $(1:-1:1)$. The quotient $\mathrm{Bl}_{3}(\mathbb{P}^{2})/T_{6}$ has three singular points, which turn out to be of types $A_{5}$, $A_{2}$ and $A_{1}$. On the minimal desingularization $M_{6}\to\mathrm{Bl}_{3}(\mathbb{P}^{2})/T_{6}$, we have a foliation induced by $\mathcal{E}_{6}$. This is Santos’s foliation $\mathcal{F}_{6}$. It has a link $C$ (in the sense of Section 1), image of $\Delta_{6}$, with $C^{2}=1$ (see [San17]).
@@
4. Chazy equations and foliations on a weighted projective plane
4.1. Chazy IV and $\mathcal{G}_{\mathrm{IV}}$
We begin by describing the relation between the reduced Chazy IV equation and Brunella’s very special foliation following [Gui12]. The Chazy IV equation (4) is given by the vector field
$W=y\frac{\partial}{\partial x}+z\frac{\partial}{\partial y}+3(xz+y^{2}-x^{2}y)\frac{\partial}{\partial z},$
which induces the foliation on $\mathbb{P}(1,2,3)$ that will be denoted by $\mathcal{G}_{\mathrm{IV}}$. The latter may also be defined by the quasihomogeneous form
@@ 4.1. Chazy IV and $\mathcal{G}_{\mathrm{IV}}$
The vector field $W$ has the quasihomogeneous first integral of degree three $B=x^{3}-3xy+z$, and the invariant surface given by the quasihomogeneous polynomial of degree six
@@
Let $\Sigma=B^{-1}(1)$. The vector field $W$ is tangent to $\Sigma$, and induces on it a foliation, that we will denote by $\widetilde{\mathcal{G}}_{\mathrm{IV}}$. There is an action of $\mathbb{Z}_{3}$ on $\mathbb{C}^{3}$ given by the restriction of the weighted homotheties (7) to the group of cubic roots of unity. It preserves $\Sigma$ along with $\widetilde{\mathcal{G}}_{\mathrm{IV}}$. The quotient of $\Sigma$ under this action is realized by the restriction to $\Sigma$ of the quotient map $\pi:\mathbb{C}^{3}\setminus\{0\}\to\mathbb{P}(1,2,3)$. The image of $\Sigma$ is the complement of the curve $B=0$ in $\mathbb{P}(1,2,3)$, and the image of $\widetilde{\mathcal{G}}_{\mathrm{IV}}$ is the restriction of $\mathcal{G}_{\mathrm{IV}}$ to this image.
On the other hand, we have the foliation $\mathcal{E}_{3}$ on $\mathbb{P}^{2}$ described in Section 3.1. There is an action of $\mathbb{Z}_{3}$ on $\mathbb{P}^{2}$ preserving $\mathcal{E}_{3}$, given by the transformation $T_{3}$ of Eq. (16); the quotient of $\mathbb{P}^{2}$ is a rational surface, and the induced foliation is Brunella’s very special foliation $\mathcal{F}_{3}$.
We will exhibit a birational map $\Phi_{3}:\mathbb{P}^{2}\dashrightarrow\Sigma$ that maps $\mathcal{E}_{3}$ to $\widetilde{\mathcal{G}}_{\mathrm{IV}}$, and that is equivariant with respect to the actions of $\mathbb{Z}_{3}$. This will establish the relation between $\mathcal{F}_{3}$ and $\mathcal{G}_{\mathrm{IV}}$ announced in Theorem 4.
@@
(23) $(\omega-1)\left(X\frac{\partial}{\partial X}-\omega Y\frac{\partial}{\partial Y}\right),$
and which is tangent to the foliation $\mathcal{E}_{3}$ of Eq. (15). Consider the rational function on $\mathbb{P}^{2}$
@@
A straightforward calculation shows that, with respect to the derivation given by $D_{3}$, $f$ is a solution to the reduced Chazy IV equation (4). (The above expression for $f$ corrects the one given in [Gui12, p. 72].)
The map $\mathbb{P}^{2}\dashrightarrow\mathbb{C}^{3}$, given by $(X:Y:1)\mapsto(f,D_{3}f,D_{3}^{2}f)$ takes values in $\Sigma$ and defines a map $\Phi_{3}:\mathbb{P}^{2}\dashrightarrow\Sigma$. Since $f$ satisfies the Chazy IV equation with respect to the derivation given by $D_{3}$, $\Phi_{3}$ maps $D_{3}$ to the restriction of $W$ to $\Sigma$, thus mapping $\mathcal{E}_{3}$ to $\widetilde{\mathcal{G}}_{\mathrm{IV}}$. The map $\Phi_{3}$ is equivariant with respect to the action of $\mathbb{Z}_{3}$ on $\mathbb{P}^{2}$ given by the transformation $T_{3}$ of Eq. (16), and to its previously described action on $\Sigma$. It is a birational one, for its inverse, when $\Sigma$ is parametrized by $(x,y)\mapsto(x,y,1-x^{3}+3xy)$, is given by
@@
In this way, Brunella’s very special foliation, given by the quotient of the foliation on $\mathbb{P}^{2}$ induced by $D_{3}$ under the action of $\mathbb{Z}_{3}$ given by Eq. (16), is birationally equivalent to the quotient of the foliation induced by $W$ on $\Sigma$, the foliation $\mathcal{G}_{\mathrm{IV}}$ on $\mathbb{P}(1,2,3)$ induced by the Chazy IV equation. Under this correspondence, the link corresponds to the curve defined by $C$.
@@
Observe that, by developing $\Phi_{3}$, we obtain an explicit birational map from $\mathbb{P}^{2}$ to $\mathbb{P}(1,2,3)$ that realizes the quotient of $\mathbb{P}^{2}$ under the action of the cyclic permutation of variables (16), under which the pull-back of (22) is the form (15) defining the foliation $\mathcal{E}_{3}$ of Section 3.1.
@@
Let us give explicit expressions in this model of $\mathcal{F}_{3}$ for the symmetries described in Section 3.1. We may conjugate the symmetries on $\mathbb{P}^{2}$ via $\Phi_{3}$, and obtain explicit birational transformations of $\Sigma$, which, on their turn, induce birational transformations of $\mathbb{P}(1,2,3)$ preserving $\mathcal{G}_{\mathrm{IV}}$. Brunella’s foliated flop, the symmetry associated to the Cremona involution (17), is the birational involution of $\mathbb{P}(1,2,3)$ given by
(24) $[x:y:z]\mapsto[2xy-z:C:(3xy-2z)C];$
the birational trivolution of $\mathbb{P}(1,2,3)$ induced by (18) is
(25) $[x:y:z]\mapsto[x^{2}-y:x^{2}y-xz+y^{2}:3x^{4}y-2x^{3}z-3x^{2}y^{2}+3xyz-2y^{3}].$
These two generate a group of birational transformations of $\mathbb{P}(1,2,3)$ preserving $\mathcal{G}_{\mathrm{IV}}$, isomorphic to the group, of order six, of permutations in three symbols $S_{3}$. This group can be promoted to a group of birational automorphisms of $\mathbb{C}^{3}$ that preserve the vector field associated to the reduced Chazy IV equation. This is the content of Theorem 6, which can be established by a direct calculation. In it, (8) corresponds to Brunella’s foliated flop, (9) to the above trivolution, and the remaining transformations to the other non-trivial elements of this group.

### v2

@@ 3.1. Brunella’s very special foliation, $\mathcal{F}_{3}$
It is tangent to the coordinate triangle $\Delta_{3}:XYZ=0$, and has three reduced non-degenerate singular points at its vertices. It is preserved by the linear automorphism of order three
@@
  • first, a blow-up $\sigma$ of the node $p$ of the link $C$ transforms it into a curve $\overline{C}$ of self-intersection 1-1;
@@ 3.2. Santos’s foliation $\mathcal{F}_{4}$
for $\mathrm{i}=\sqrt{-1}$. It is tangent to the cycle $\Delta_{4}$ of four lines formed by $\mathbb{P}^{1}\times\{0\}$, $\mathbb{P}^{1}\times\{\infty\},$ $\{0\}\times{\mathbb{P}}^{1}$, and $\{\infty\}\times\mathbb{P}^{1}$, and has reduced non-degenerate singularities at its vertices. The order-four automorphism $T_{4}$ of $\mathbb{P}^{1}\times\mathbb{P}^{1}$
@@
permutes cyclically the four lines of $\Delta_{4}$, and preserves the foliation $\mathcal{E}_{4}$. It acts freely in a neighborhood of $\Delta_{4}$. The transformation $T_{4}$ has two fixed points, $(1:1,1:1)$ and (1:1,1:1)(-1:1,-1:1), at which the eigenvalues of the derivative of $T_{4}$ are $\mathrm{i}$ and $\mathrm{i}$. It also has an orbit of length two, formed by (1:1,1:1)(-1:1,1:1) and (1:1,1:1)(1:1,-1:1), at which the derivative of $T^{2}$ has twice the eigenvalue 1-1. The variety $(\mathbb{P}^{1}\times\mathbb{P}^{1})/T_{4}$ has four singular points, two of type $A_{3}$ and one of type $A_{1}$. The minimal desingularization $M_{4}\to(\mathbb{P}^{1}\times\mathbb{P}^{1})/T_{4}$ is endowed with a foliation, quotient of $\mathcal{E}_{4}$. This is Santos’s foliation $\mathcal{F}_{4}$; it has a link $C$ (in the sense of Section 1), image of $\Delta_{4}$, with $C^{2}=2$ (see [San17]).
@@ 3.3. Santos’s foliation $\mathcal{F}_{6}$
Let us consider again the foliation $\mathcal{E}_{3}$ on $\mathbb{P}^{2}$ given in Section 3.1 by Eq. (15). Let $\sigma:\mathrm{Bl}_{3}(\mathbb{P}^{2})\to\mathbb{P}^{2}$ be the composition of the blowing-ups at the three vertices of the coordinate triangle: $(0:0:1)$, $(0:1:0)$, and $(1:0:0)$. Consider the cycle $\Delta_{6}$ of six (1)(-1)-curves on $\mathrm{Bl}_{3}(\mathbb{P}^{2})$ formed by the transform of the coordinate triangle of $\mathbb{P}^{2}$ by $\sigma$, and by the three exceptional lines of $Q$. Denote by $\mathcal{E}_{6}$ the transformed foliation of $\mathcal{E}_{3}$ by $Q$; it is tangent to $\Delta_{6}$, and has reduced non-degenerate singularities at its six vertices. The linear transformation $T_{3}$ of $\mathbb{P}^{2}$ of Eq. (16) commutes with the Cremona involution $Q$ in Eq. (17), and $T_{3}\circ Q$ is a birational automorphism of order six of $\mathbb{P}^{2}$ preserving $\mathcal{L}$. It induces a biholomorphism $T_{6}$ of $\mathrm{Bl}_{3}(\mathbb{P}^{2})$, which permutes cyclically the six rational curves of $\Delta_{6}$, and preserves the foliation $\mathcal{E}_{6}$. The automorphism $T_{6}$ has a fixed point coming from $(1:1:1)$, the common fixed point of $T_{3}$ and $Q$; an orbit of order two coming from the other fixed points of $T_{3}$, $(1:\omega:\omega^{2})$ and $(1:\omega^{2}:\omega)$; and an orbit of order three, formed by the other fixed points of $Q$, (1:1:1)(1:1:-1), (1:1:1)(-1:1:1) and (1:1:1)(1:-1:1). The quotient $\mathrm{Bl}_{3}(\mathbb{P}^{2})/T_{6}$ has three singular points, which turn out to be of types $A_{5}$, $A_{2}$ and $A_{1}$. On the minimal desingularization $M_{6}\to\mathrm{Bl}_{3}(\mathbb{P}^{2})/T_{6}$, we have a foliation induced by $\mathcal{E}_{6}$: this is Santos’s foliation $\mathcal{F}_{6}$. It has a link $C$ (in the sense of Section 1), image of $\Delta_{6}$, with $C^{2}=1$ (see [San17]).
Remark 13.
@@
4. The Chazy equations and the special quotient foliations
In this section we will study the relation between the special quotient foliations of the previous section and the Chazy IV, V and VI equations, and prove Theorem 4. The plane models for the special quotient foliations of Theorems 2 and 3, and an alternative proof of Theorem 1, will be deduced from this. The symmetries of the special quotient foliations presented in the previous section will give symmetries for the Chazy equations, yielding Theorems 6 and 7. Also, the relation between $\mathcal{F}_{3}$ and $\mathcal{F}_{6}$ of Remark 13 will give, through the relations here explored, the result stated in Theorem 8.
4.1. The Chazy IV equation and Brunella’s foliation $\mathcal{F}_{3}$
In this section we will establish the part of Theorem 4 concerning the birational equivalence of Brunella’s very special foliation $\mathcal{F}_{3}$ and the foliation on $\mathbb{P}(1,2,3)$ induced by the Chazy IV equation. As a consequence of this equivalence, we will obtain another proof of Theorem 1. Also, through it, the symmetries of $\mathcal{F}_{3}$ described in the Section 3.1 will produce the symmetries of the Chazy IV equation featured in Theorem 6.
4.1.1. Chazy IV and $\mathcal{F}_{3}$
The Chazy IV equation (4) is given by the vector field on $\mathbb{C}^{3}$
$W=y\frac{\partial}{\partial x}+z\frac{\partial}{\partial y}+3(xz+y^{2}-x^{2}y)\frac{\partial}{\partial z}.$
It is quasihomogeneous with respect to the weights $1$, 22 and $3$ for $x$, $y$ and $z$, respectively. The vector field induces a foliation on $\mathbb{P}(1,2,3)$ that we will denote by $\mathcal{G}_{\mathrm{IV}}$. The latter may also be defined by the quasihomogeneous form
@@ 4.1.1. Chazy IV and $\mathcal{F}_{3}$
The discussion that follows is based on [Gui12]. The vector field $W$ has the quasihomogeneous first integral of degree three $B=x^{3}-3xy+z$, and the invariant surface given by the quasihomogeneous polynomial of degree six
@@
Let $\Sigma=B^{-1}(1)$. The vector field $W$ is tangent to $\Sigma$, and induces on it a foliation, that we will denote by $\widetilde{\mathcal{G}}_{\mathrm{IV}}$. Let $\omega$ be a primitive cubic root of unity. There is an action of $\mathbb{Z}_{3}$ on $\mathbb{C}^{3}$ given by the restriction of the weighted homotheties (7) to the cubic roots of unity, generated by
(23) $(x,y,z)\mapsto(\omega x,\omega^{2}y,z).$
This action preserves $\Sigma$, and multiplies $W$ by $\omega$, thus preserving $\widetilde{\mathcal{G}}_{\mathrm{IV}}$. The quotient of $\Sigma$ under this action is realized by the restriction to $\Sigma$ of the quotient map $\pi:\mathbb{C}^{3}\setminus\{0\}\to\mathbb{P}(1,2,3)$. The image of $\Sigma$ is the complement of the curve $B=0$ in $\mathbb{P}(1,2,3)$, and the image of $\widetilde{\mathcal{G}}_{\mathrm{IV}}$ is the restriction of $\mathcal{G}_{\mathrm{IV}}$ to this image.
For the foliation $\mathcal{E}_{3}$ on $\mathbb{P}^{2}$ described in Section 3.1, there is an action of $\mathbb{Z}_{3}$ on $\mathbb{P}^{2}$ preserving it, given by the transformation $T_{3}$ of Eq. (16); the quotient of $\mathbb{P}^{2}$ under this action is a rational surface, and the induced foliation is Brunella’s very special foliation $\mathcal{F}_{3}$.
In order to establish the birational equivalence between $\mathcal{F}_{3}$ and $\mathcal{G}_{\mathrm{IV}}$ stated in Theorem 4, we will exhibit a birational map $\Phi_{3}:\mathbb{P}^{2}\dashrightarrow\Sigma$ that maps $\mathcal{E}_{3}$ to $\widetilde{\mathcal{G}}_{\mathrm{IV}}$, and that is equivariant with respect to the actions of $\mathbb{Z}_{3}$.
@@
(24) $(\omega-1)\left(X\frac{\partial}{\partial X}-\omega Y\frac{\partial}{\partial Y}\right).$
It is tangent to the foliation $\mathcal{E}_{3}$ of Eq. (15). Consider the rational function on $\mathbb{P}^{2}$
@@
A lengthy but nevertheless straightforward calculation (which we omit) shows that, with respect to the derivation given by $D_{3}$, $f$ is a solution to the reduced Chazy IV equation (4). (The above expression for $f$ corrects the one given in [Gui12, p. 72].) The map $\mathbb{P}^{2}\dashrightarrow\mathbb{C}^{3}$, given by $(X:Y:1)\mapsto(f,D_{3}f,D_{3}^{2}f)$ takes thus values in a level set of $H$; this level set is $\Sigma$. We thus have a map $\Phi_{3}:\mathbb{P}^{2}\dashrightarrow\Sigma$. Since $f$ satisfies the Chazy IV equation with respect to the derivation given by $D_{3}$, $\Phi_{3}$ maps $D_{3}$ to the restriction of $W$ to $\Sigma$, thus mapping $\mathcal{E}_{3}$ to $\widetilde{\mathcal{G}}_{\mathrm{IV}}$. The map $\Phi_{3}$ is equivariant with respect to the action of $\mathbb{Z}_{3}$ on $\mathbb{P}^{2}$ given by the transformation $T_{3}$ of Eq. (16), and to the action of Eq. (23) on $\Sigma$. The map $\Phi_{3}$ is also birational: its birational inverse is, when $\Sigma$ is parametrized by $(x,y)\mapsto(x,y,1-x^{3}+3xy)$, given by
@@
This establishes the birational equivalence between $\mathcal{F}_{3}$ and $\mathcal{G}_{\mathrm{IV}}$ stated in Theorem 4.
@@
The quotient of $\mathbb{P}^{2}$ under the action of the cyclic permutation of variables (16) is realized by the explicit rational map from $\mathbb{P}^{2}$ to $\mathbb{P}(1,2,3)$ given by $\pi\circ\Phi_{3}$. Under this map, the pull-back of the form of Eq. (22) is the form of Eq. (15), the one defining the foliation $\mathcal{E}_{3}$ of Section 3.1.
4.1.2. Birational symmetries of Chazy IV and $\mathcal{G}_{\mathrm{IV}}$
Let us study the symmetries of $\mathcal{G}_{\mathrm{IV}}$ induced by the symmetries of $\mathcal{F}_{3}$ described in Section 3.1. Theorem 6 will be a consequence of this.
@@ 4.1.2. Birational symmetries of Chazy IV and $\mathcal{G}_{\mathrm{IV}}$
The linear symmetries of $\mathbb{P}^{2}$ that induce the symmetries of $\mathcal{F}_{3}$ discussed in Section 3.1 may be conjugated by the previous transformation $\Phi_{3}$ and its inverse, in order to obtain explicit birational transformations of $\Sigma$, which, on their turn, induce birational transformations of $\mathbb{P}(1,2,3)$ preserving $\mathcal{G}_{\mathrm{IV}}$. In this way, Brunella’s foliated flop, the symmetry associated to the Cremona involution (17), is found to be the birational involution of $\mathbb{P}(1,2,3)$ given by
(25) $x:y:z]\mapsto[2xy-z:C:(3xy-2z)C],$
while the birational trivolution of $\mathbb{P}(1,2,3)$ induced by (18) is
(26) $x:y:z]\mapsto[x^{2}-y:x^{2}y-xz+y^{2}:3x^{4}y-2x^{3}z-3x^{2}y^{2}+3xyz-2y^{3}].$
These two generate a group of birational transformations of $\mathbb{P}(1,2,3)$ preserving $\mathcal{G}_{\mathrm{IV}}$, isomorphic to the group of permutations in three symbols $S_{3}$.
These symmetries are behind Theorem 6, which can be established by a direct calculation, one whose inclusion here would be of little interest. We will nevertheless explain how the expressions in Theorem 6 were obtained, as well as their relation with the above symmetries.
The transformation (8) of Theorem 6 is a lift of the involution (25) to $\mathbb{C}^{3}$. In fact, when extending (8) to $x^{\prime}$ and $x^{\prime\prime}$ based on the way in which $x$ is transformed, and on the differential equation solved by $x$, we obtain
$(x,x^{\prime},x^{\prime\prime})\mapsto\left(\frac{2xx^{\prime}-x^{\prime\prime}}{x^{\prime}},\frac{C(x,x^{\prime},x^{\prime\prime})}{(x^{\prime})^{2}},\frac{(3xx^{\prime}-2x^{\prime\prime})C(x,x^{\prime},x^{\prime\prime})}{(x^{\prime})^{3}}\right),$
which induces the transformation (25) on $\mathbb{P}(1,2,3)$. In a completely analogous way, the transformation (9) of Theorem 6 is related to the trivolution (26). The group of birational automorphisms of $\mathbb{P}(1,2,3)$ generated by (25) and (26) can be thus promoted to a group of birational automorphisms of $\mathbb{C}^{3}$ that preserve the vector field associated to the reduced Chazy IV equation.
Let us explain how the transformation (8) was obtained. The natural lift of (25) as a polynomial self-map of $\mathbb{C}^{3}$ has the form $(x,y,z)\mapsto(P,Q,R)$, for $P$, $Q$ and $R$ quasihomogeneous polynomials of degrees three, six and nine, respectively. For every polynomial $S$, the transformation
(27) $(x,y,z)\mapsto\left(\frac{P}{S},\frac{Q}{S^{2}},\frac{R}{S^{3}}\right)$
gives a rational self-map of $\mathbb{C}^{3}$ that induces the transformation (25) on $\mathbb{P}(1,2,3)$. If we want such a transformation to be a symmetry of the Chazy IV equation, then, to begin with, $S$ must be quasihomogeneous of degree two, for only in that case the components in the right-hand side of (27) have degrees one, two and three, as $x$, $y$ and $z$ do: if a transformation of the form (27) preserves the Chazy IV equation, we should have that $S=\alpha x^{2}+\beta y$ for some $\alpha,\beta\in\mathbb{C}$. If we impose the first necessary condition $W(P/S)=Q/S^{2}$ on such $S$, the only possibility is found to be $S=y$. In this way we find the transformation (8) of Theorem 6. The transformation (9) is obtained in an analogous way; the remaining transformations are obtained by composing these two.
4.1.3. Description of $\mathcal{G}_{\mathrm{IV}}$